OpenStudy (anonymous):

Factor completely p4 – 16 last question i have i need help

OpenStudy (anonymous):

help please

OpenStudy (earthcitizen):

you posted this a second time ?

OpenStudy (earthcitizen):

http://openstudy.com/#/updates/4efe3ffce4b01ad20b5302a6

OpenStudy (asnaseer):

@TuringTest has already given you a hint at how to solve this

OpenStudy (anonymous):

yeah cuz i have no clue how to do this

OpenStudy (asnaseer):

do you understand this factorization: \[a^2-b^2=(a+b)(a-b)\]

OpenStudy (anonymous):

is a and b x and y?

OpenStudy (anonymous):

like what would a and b be

OpenStudy (wasiqss):

(x^2-4)(x^2+4)

OpenStudy (asnaseer):

in your case:\[ a^2 = p^4\]\[b^2=16\]

OpenStudy (wasiqss):

(p^2-4)(p^2+4)

OpenStudy (anonymous):

okay

OpenStudy (wasiqss):

done it

OpenStudy (turingtest):

it can be factored again... listen to what asnaseer is saying about the form in general

OpenStudy (anonymous):

this is what is confusing me; A.(p2 – 4)2 B.(p2 + 2)(p2 + 8) C. (p2 + 4)(p + 4)(p – 4) D. (p2 + 4)(p + 2)(p – 2) at wasiqss your wrong it is something different cuz what u did i would of done to but its wrong

OpenStudy (asnaseer):

look at what you now have left:\[p^4-16=(p^2-4)(p^2+4)\] notice the "form" of (p^2-4)

OpenStudy (anonymous):

okay i have that written down

OpenStudy (asnaseer):

see if you can apply the same rule again to this

OpenStudy (anonymous):

so i take the p^2-4 and apply it to the difference of squares?

OpenStudy (asnaseer):

so you know:\[a^2-b^2=(a+b)(a-b)\] now:\[a^2=p^2\]\[b^2=4\]

OpenStudy (asnaseer):

yes - I think you are getting the hang of it now - well done

OpenStudy (anonymous):

i got p^2-4^2=(p+4)(p-4)

OpenStudy (asnaseer):

bingo! - you're a genius!

OpenStudy (asnaseer):

oops - sorry candycane you should have written:\[p^2-4=(p-2)(p+2)\]

OpenStudy (asnaseer):

remember \(b^2=4\), so \(b=2\)

OpenStudy (asnaseer):

so now if you put all the components together, what do you get?

OpenStudy (anonymous):

(p2 + 4)(p + 2)(p – 2)

OpenStudy (turingtest):

yay! this time definitely right :)

OpenStudy (asnaseer):

thats it - you have the right answer - well done again. the trick is in spotting when to use the difference of two squares factorization.

OpenStudy (anonymous):

thank you soooo much! i definatly get it now

OpenStudy (asnaseer):

yw